Ask
your own question, for FREE!
Mathematics
16 Online
Brian has a bag that contains 14 red marbles and 12 yellow marbles. He selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Brian selects a red marble and then a yellow marble? Round your answer to the nearest percent. P(red and yellow)
Still Need Help?
Join the QuestionCove community and study together with friends!
26%
46%
The correct answer is 26% because 14/26 * 12/25 = 0.25846154.
26%
probability to get a red marble is 14/26 whereas the probability to get an yellow marble is 12/25 I supposed to get a red marble for first. So the requested probability is: \[\huge p = \frac{{14}}{{26}} \times \frac{{12}}{{25}} \times 100 = ...?\]
Still Need Help?
Join the QuestionCove community and study together with friends!
so 26%
yes! We get \(25.8 \%\) which can be rounded to \(26\%\)
k
:)
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
15 hours ago
0 Replies
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal